OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation −9 + 8v = 7 + 4v is given below. Choose the step that is missing from the solution. −9 + 8v = 7 + 4v 4v = 16 v = 4 WHICH ANSWER IS IT? : 8v = −2 + 4v −9 + 12v = 7 8v = 16 + 4v −2 + 8v = 4v

OpenStudy (anonymous):

The first step is to add 9 to both sides, so you have (-9 + 9) + 8v = (7+9)+ 4v This gives you 8v = 16 + 4v = the 3rd step listed above.